40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 29 03:16:21 2016, ericq, Update, LSC, DRFPMI Locked Once More 2016-03-29_calibdarm.pdf2016-03-29_DRFPMI_errctrl.pdfDRFPMI_DTT.zip
    Reply  Wed Mar 30 16:40:24 2016, ericq, Update, LSC, 2016 vs 2010 2016_vs_2010.png
    Reply  Mon Apr 11 16:06:30 2016, ericq, Update, LSC, DRFPMI Data Archived 
Message ID: 12053     Entry time: Tue Mar 29 03:16:21 2016     Reply to this: 12055   12069
Author: ericq 
Type: Update 
Category: LSC 
Subject: DRFPMI Locked Once More 

[ericq, Gautam]

Three RF-only locks longer than a minute tonight, out of 5 total attempts. 

Last week, I determined that the beam spot on the RF POP PD is too large. This still needs to be fixed. I updated the ASS model to use REFLDC as a PRCL dither error signal; it works. 

There seems to be some excess angular motion of ETMY tonight. This is evident in the oplev spectra (as compared to ETMX), and the GTRY camera, and even the retroreflected beam from a misalgined ETMY on the ITMY face when the PRC is carrier locked.

Gautam and I mostly focused on setting up the CAL-DARM_CINV block to produce this (mostly) calibrated spectrum starting from GPS 1143274087. [Darm on unwhitened AS55, DRMI on 3F, one CARM boost]

Here are the control and error signal spectra:

[DTT files attached]

Note to self: archive some of this data 

Attachment 1: 2016-03-29_calibdarm.pdf  47 kB  | Hide | Hide all | Show all
2016-03-29_calibdarm.pdf
Attachment 2: 2016-03-29_DRFPMI_errctrl.pdf  128 kB  | Show | Hide all | Show all
Attachment 3: DRFPMI_DTT.zip  2.517 MB
ELOG V3.1.3-